Two sides of a pentagon have length (3x-5) and three sides of length (2x+1). If the perimeter is 89, what is the length of the longest side?
Answer by robertb(5830)

2(3x-5) + 3(2x+1) = 89,
6x -10 + 6x+3 = 89,
12x - 7 = 89,
12x = 96,
x = 8. Each of the two sides has length 19, and each of the three sides has length 17. Now you know the answer...